OpenStudy (anonymous):

Simplify: (-4)/(x^2-16)+x/(2x+8)-1/4

OpenStudy (agreene):

\[\frac{-4}{x^2-16}+\frac{x}{2x+8}-\frac14\] \[\frac{-4}{(x-4)(x+4)}+\frac{x}{2(x+4)}-\frac14\] would be where to start.

OpenStudy (anonymous):

alright thanks.